A patient is to receive 1 unit of PRBCs over 4 hours. The label on the bag states: 300 ml. What will the flow rate be?
Answer #1

Ans) Order: A patient is to receive 1 unit of PRBCs

Time: 4 hours.

The label on the bag states: 300 ml

Formula: Flow rate ( ml/min ): Volume in ml × drop factor / time in mins.

= 300 × 10 /4 × 60

= 12.5 ml/ min

If transfusion via Infusion pump:

Flow rate ( ml/hr ) = Volume in ml / time in hrs

= 300/4 = 75 ml/hr.